OpenStudy (anonymous):

How do you find the inverse of f(x)=x^2-2x, x less than or equal to 1?? Can you simplify the inverse and state the domain of f^-1??

OpenStudy (anonymous):

What is the problem exactly? With inverse do you mean 1/f(x)? (or f(x)^-1 if you prefer that notation. can't you simply\[\frac{ 1 }{ x^2-2x }\] and realize that you can't divide by 0 so not all values of x will be allowed?

OpenStudy (anonymous):

i meant f(x) @zimmah

OpenStudy (anonymous):

but what do you need to solve?

OpenStudy (anonymous):

the domain of the inverse of f(x) is any x that does not violate the rule of division by 0 if that's what you mean.

OpenStudy (anonymous):

I dont know, thats all this problem says @zimmah

OpenStudy (anonymous):

Is that literally how they formulated it?

OpenStudy (anonymous):

yes. @zimmah

OpenStudy (anonymous):

In that case my interpretation of the answer would be \[f(x)^{-1}=\frac{ 1 }{ x^2-2x } \in x < 0\] not that x=0 is not included because it is undefined at that point.